OpenStudy (anonymous):

If c is the measure of the hypotenuse of a right triangle, find each missing measure. If necessary, round to the nearest hundredth. a = 60, b = ?, c = 100 Select one: a. 160 b. 40 c. 116.62 d. 80.00

OpenStudy (zzr0ck3r):

 a^2+b^2 = c^2

OpenStudy (zzr0ck3r):

\[60^2+b^2 = 100^2\] can you solve for b?

OpenStudy (anonymous):

b=80,-80 is this right?

OpenStudy (zzr0ck3r):

correct

OpenStudy (zzr0ck3r):

but triangles don't have negative side length so take 80

OpenStudy (anonymous):

ok, thank you.